Sue bought a blouse on salr that was 30% less than the original price. The original price was $15 more than the sale price.what was the original price? Show work

Answer by abhiverma27(3)

Let the original price of the blouse be X
Therefore 30% less of original price means 30% of X i.e. 30/100*X = 3/10*X
now original price was $15 more than the sales price, which means the discount which is given is equal to $15
so 3/10*X = 15
X = 15*10/3
X = 50
that's your answer. Good luck.